A business firms sells a good at the price of Rs 450.The firm has decided to reduce the
price of good to Rs 350.Consequently, the quantity demanded for the good rose from
25,000 units to 35,000 units. Calculate the price elasticity of demand.
1
Expert's answer
2020-12-17T07:46:10-0500

"Price elasticity o demand = \frac{percentage change in demand}{ percentage change in price}"


Change in demand = D2 - D1 = 35,000- 25,000 = 10,000


Mid point = "\frac{25000+35000}{2} = 30,000"


percentage change in demand = "\frac{10000}{30000} \times100= 33\frac{1}{3} %"


change in price = P2 - P1 = 350 - 450 = -100


Mid point = "\frac{350 + 450}{2} = 400"


percentage change in price = "\frac{-100}{400} \times100 = -25"


So,

Price elasticity of demand = "\frac{33.33}{25} = - 1\frac{1}{3}"